OpenStudy (anonymous):

y-intercept form of this equation? 8y^2=x(x-4)^2

OpenStudy (confluxepic):

Substitute x=0 and solve for y.

OpenStudy (confluxepic):

Then you put the y-value as (0,y) and that would be your y-intercept.

OpenStudy (anonymous):

oh no I've made a mistake

OpenStudy (anonymous):

I have to solve for y

OpenStudy (campbell_st):

let x = 0 so its 8y^2 = 0(0 -4)^2 so 8y^2 = 0 should be easy to solve now